OpenStudy (anonymous):

subtitude and then simplify: 30 degrees = 1/2 simplify: 5 sin^2 30 degrees

OpenStudy (anonymous):

I'm not sure which to subtitude for

OpenStudy (p0sitr0n):

5 *0.5^2 = 0.25*5 = 1.25 ?

OpenStudy (p0sitr0n):

substitute sin 30 by 1/2

OpenStudy (anonymous):

so i would plug in Sin 1/2 ?

OpenStudy (p0sitr0n):

it will be: 5sin^2 30 5 (0.5)^2

OpenStudy (anonymous):

ohh ok i thought you said after u get 1.25 then subtitude for 1/2 again, i was a little confuse there lol

OpenStudy (anonymous):

so 1.25 is the answer ?

OpenStudy (p0sitr0n):

yep